### v4

# Video LLMs for Temporal Reasoning in Long Videos

Abstract
This paper introduces TemporalVLM, a video large language model (video LLM) capable of effective temporal reasoning and fine-grained understanding in long videos. At the core, our approach includes a visual encoder for mapping a long-term input video into features which are time-aware and contain both local and global cues. In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps and fused across overlapping temporal windows into time-sensitive local features. Next, the local features are passed through a bidirectional long short-term memory (BiLSTM) module for global feature aggregation. The extracted time-aware and multi-level features are important for accurate temporal reasoning and fine-grained understanding in long videos. Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, which consists of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation. Finally, extensive experiments on datasets of long videos, including TimeIT and IndustryASM, show that TemporalVLM achieves superior performance than previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. To the best of our knowledge, our work is the first to incorporate LSTMs into video LLMs.
Figure 1: Previous video large language models (a-c) usually are not time-sensitive (a, b), consider an input video as a single clip (a, c), and apply pooling operation (a, b) or query aggregation (c) for aggregating global semantic information. In contrast, our model (d) includes a time-aware clip encoder, which extracts time-aware fine-grained cues from short-term clips sampled from a long-term input video, and a bidirectional long short-term memory module, which captures long-range temporal dependencies across multiple clips. Our extracted time-aware and multi-level features are crucial for temporal reasoning in long videos.
1 Introduction
00footnotetext: indicates joint first author.
{fawad,umer,hamza,zeeshan,huy}@retrocausal.ai.
Video temporal reasoning represents the process of reasoning about time and its passage in videos, with a focus on how events or actions happen and relate to each other in terms of time. Several video understanding applications require the ability of temporal reasoning. Examples include dense video captioning [46, 60, 53], which attempts to detect the timestamps and generate detailed descriptions of all the events or actions in an input video, and temporal action segmentation [23, 54, 2, 26], which aims to associate each frame of a long video capturing a complex activity with one of the action/sub-activity classes. Considerable efforts [53, 27, 30, 26] have been invested in developing separate models for solving individual tasks. These models usually have different architectures. Therefore, it is favorable to design a unified model for handling various tasks.
The past few years have witnessed the impressive instruction following ability and remarkable comprehension and generation capability of large language models (LLMs) [1, 4, 41, 31, 42, 43], which have emerged as a universal agent for performing various tasks. Video LLMs [22, 24, 28, 29, 57, 16, 39] which incorporate video encoders with LLMs for video understanding tasks have been introduced. These methods often represent a video by a fixed number of visual tokens, leading to reduced performance with long videos, and encode video frames and their timestamps separately and hence struggle to perform temporal reasoning tasks. Recently, a few works [33, 13, 14, 36] have developed video LLMs with temporal reasoning abilities, e.g., TimeChat [36] proposes to vary the number of final video tokens based on the input video length and jointly encode video frames and their timestamps. The above methods usually treat the entire video as a single clip [33, 13, 14, 36] and aggregate video tokens via pooling operation [14] and query aggregation [36], struggling to capture fine-grained information, especially for long videos.
We propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos. Fig. 1 shows the differences in model architectures of TemporalVLM and previous works. Our model includes two key architectural contributions for temporal reasoning in long videos. Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting fused time-aware local features from each clip. Secondly, we adopt a bidirectional long short-term memory (BiLSTM) module which takes all the local features as inputs and computes global features from multiple clips. As a consequence, our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos. Moreover, to further evaluate our model, we present IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation benchmarking and time and motion studies. Our IndustryASM dataset comprises of $4803$ videos with an average video duration of $119$ seconds. It covers in total $52$ diverse industry assembly tasks and includes timestamp and action labels. We convert the labels into chat samples with manually written instructions. To the best of our knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers for temporal action segmentation evaluation and time and motion analyses. Lastly, extensive experiments on datasets of long videos, i.e., TimeIT [36] and IndustryASM, demonstrate that our TemporalVLM approach achieves superior results over previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
In summary, our contributions include:
  • We develop TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos. Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into fused time-aware local features. The local features are then fed to a BiLSTM module to extract global features. Therefore, our features not only are time-sensitive but also capture effective local and global information.
  • We present IndustryASM, a large-scale long video dataset of manufacturing assembly procedures. Our videos are recorded on factory floors with actions and timestamps labeled by industrial engineers for time and motion analyses and temporal action segmentation evaluation. We will make our dataset available publicly.
  • Our evaluations show that TemporalVLM outperforms prior works across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to integrate LSTMs into video LLMs.
2 Related Work
Large Language Models. Over the past few years, LLMs [1, 4, 41, 31, 42, 43] have revolutionized the field of natural language processing. Initial methods employ encoder-decoder architectures [17, 35], whereas decoder-only models, e.g., GPT [34], demonstrate impressive performance and generalization in recent years. In addition, models with billions of parameters, e.g., GPT-4 [1], have been developed to push the boundary further, while instruction tuning techniques [32, 47, 48] have been widely used to enable models to generate contextually relevant responses. Here, we focus on video LLMs, which take videos as additional inputs and explore the use of LLMs for video understanding tasks.
Video Large Language Models. Motivated by the impressive reasoning and generalization abilities of LLMs, video LLMs [22, 24, 28, 29, 57, 16, 39] have been developed for video understanding tasks. They typically include three key components: i) a pre-trained visual encoder (e.g., ViT [38]) to extract visual features, ii) a projection layer to map visual features into the text latent space of LLMs, and iii) a pre-trained LLM (e.g., LLaMA [42]) for generating responses. They mostly differ in their visual encoder, i.e., how visual features are extracted and aggregated from an input video. For example, VideoChat [22] extracts frame features via a visual transformer [38] and employs a query transformer (Q-Former) [21] to aggregate frame features into video features, while a video Q-Former is further included for temporal modeling in Video-LLaMA [57]. The above methods usually map the video into a fixed number of visual tokens, yielding degrading performance with long videos, while encoding video frames and their timestamps separately and hence struggling with temporal reasoning tasks. To address these drawbacks, methods with temporal reasoning capabilities have been introduced, e.g., TimeChat [36] presents a sliding video Q-Former to handle various video lengths and a time-aware frame encoder to jointly encode video frames and their timestamps. These methods often consider the video as a single clip [28, 29, 22, 57] and aggregate video tokens via pooling operation [28, 29] and query aggregation [22, 57], overlooking fine-grained details. In this work, we divide the video into multiple clips and propose a time-aware clip encoder for capturing local features from each clip. Also, we introduce a BiLSTM module for aggregating global features over multiple clips. To our best knowledge, our work is the first to exploit LSTMs for video LLMs.
Video Temporal Reasoning. Temporal reasoning plays an important role in many video understanding tasks, e.g., dense video captioning [46, 60, 53], temporal video grounding [49, 27], video highlight detection [20, 30], and temporal action segmentation [23, 54, 2, 26]. Previous works [53, 27, 30, 26] often focus on designing separate models for tackling individual tasks, while LLM-based models [33, 13, 14, 36] capable of handling multiple tasks have emerged recently. Our TemporalVLM model belongs to the second group. Also, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion analyses.
Long Video Understanding. There exist a few challenges in long video understanding such as complex spatial-temporal relationships and redundant information. Several methods have been developed for long video understanding, including efficient architectures [6, 18], temporal pooling/aggregation [37, 51], and dynamic clip selection [9, 10]. For vision-language understanding tasks, methods based on temporal alignment [11, 3] have been introduced, while memory techniques [52, 58] are often used in dense video prediction tasks. Recently, Weng et al. [50] propose LongVLM for long video understanding. It divides the video into clips and employs a merging module for extracting local features and pooling operation for computing global features from multiple clips. It does not utilize timestamps, which are crucial for temporal reasoning. In contrast, our TemporalVLM model explicitly utilizes timestamps via a time-aware clip encoder, yielding time-aware features. Moreover, we employ a learnable BiLSTM module for aggregating global features from multiple clips.
3 TemporalVLM: A Video LLM for Temporal Reasoning in Long Videos
Figure 2: Given a long-term input video with timestamps, we divide it into short-term clips with timestamps. The time-aware clip encoder is employed on each clip along with its timestamps for extracting fused time-aware local features, which capture local fine-grained cues within each clip. Next, the bidirectional long short-term memory module is adopted for computing global features, which capture global temporal relationships across multiple clips. Lastly, the large language model takes video and text tokens as inputs and generates responses.
Below we describe our first contribution, TemporalVLM, a video LLM for effective temporal reasoning in long videos. Two novel components are proposed in TemporalVLM: i) a time-aware clip encoder for extracting time-aware local features from a short-term clip, and ii) a BiLSTM module for computing global features from the local features. We provide an overview of our TemporalVLM model in Fig. 2.
3.1 Time-Aware Clip Encoder
Video LLMs often consider a long-term input video as a single clip and uniformly sample a fixed number of frames $N^{v}_{f}$ from the video, e.g., $N^{v}_{f}=96$ frames in TimeChat [36] and $N^{v}_{f}=100$ frames in VTimeLLM [13]. Moreover, prior works usually use query aggregation, e.g., TimeChat [36], and pooling operation, e.g., Video-ChatGPT [29], to aggregate video tokens. Thus, they often perform well with short-term videos but miss out important fine-grained information for long-term videos, leading to degrading performance.
We propose a time-aware clip encoder (i.e., blue block in Fig. 2) for extracting local fine-grained cues within a clip. Given a long-term input video, we first decompose it into $C$ short-term clips and uniformly sample a fixed number of frames $N^{c}_{f}$ from each clip. Here, we set $C=6$ clips and $N^{c}_{f}=96$ frames per clip. The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding fused time-aware local features. The time-aware frame encoder uses a pre-trained image encoder [40] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [5], yielding time-aware frame features $\mathbf{f}_{t}$. These frame features $\mathbf{f}_{t}$ are then fed to a video Q-Former  [36] across overlapping windows $\mathbf{W}_{i}$ of size $q$ = 32 and overlap $o$ = 16 for temporal modeling. The window-wise outputs $\mathbf{V}_{i}$ are then concatenated into $\mathbf{S}$ :
[…]
$\mathbf{S}$ $=[\mathbf{V}_{1},\mathbf{V}_{2},\dots,\mathbf{V}_{n}],$ (2)
with $n$ denoting the number of overlapping windows within a short-term clip. We propose a fusion module to align diverse temporal cues from redundant boundary tokens in $\mathbf{S}$. Specifically, we apply multiheaded self-attention on $\mathbf{S}$, yielding $\mathbf{C}$ which fuses local contexts across multiple windows into a single context-aware embedding as:
$\mathbf{C}^{(h)}$ $=\text{SoftMax}\left(\frac{\mathbf{Q}^{(h)}{\mathbf{K}^{(h)}}}{\sqrt{d}}\right)\mathbf{V}^{(h)},$ (3)
$\mathbf{C}$ $=[\mathbf{C}^{(1)},\mathbf{C}^{(2)},\dots,\mathbf{C}^{(H)}]\mathbf{W}^{O}$ (4)
We first project $\mathbf{S}$ into $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ for each attention head $h$. At each head $h$, $\mathbf{C}^{(h)}$ represents the dot product attention between queries $\mathbf{Q}^{(h)}$ and keys $\mathbf{K}^{(h)}$ scaled over dimension $d$ followed by softmax weighted aggregation over $\mathbf{V}^{(h)}$. $\mathbf{C}$ contains the aggregated outputs from all heads with final projection $\mathbf{W}^{O}$ applied. Using overlapping windows leads to spatially redundant time-aware tokens in $\mathbf{S}$, each with different local window context. Through Eq. 4, each redundant time-aware token in $\mathbf{S}$ is enabled to compare and weigh its contextually dissimilar variant, aligning them into a fused temporal representation. This allows $\mathbf{C}$ to yield rich time-aware clip level information by leveraging diverse temporal views from across overlapping windows. Recent works, e.g., [50], leverage clip sampler and clip encoder for long video understanding but their encoders are not timeaware and hence are not effective for temporal reasoning.
3.2 Bidirectional Long Short-Term Memory
Our time-sensitive clip encoder is applied on each short-term clip to obtain time-aware local features, which capture useful fine-grained cues for temporal reasoning within the short-term clip. However, they are not effective for temporal reasoning over the long-term video, which requires the ability to capture long-range temporal dependencies.
We introduce a BiLSTM module (i.e., green block in Fig. 2) for computing global features across multiple clips. We first concatenate time-aware local features extracted from clips in the temporal order which the clips appear in the video. We then pass the sequence of fused local features to a BiLSTM module for aggregating global features. Our BiLSTM module follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward LSTM) and another processes the sequence in the reverse order (backward LSTM) as:
$\mathbf{h}^{f}_{t}=\text{LSTM}(\mathbf{h}^{f}_{t-1},\mathbf{c}_{t}),$ (5)
$\mathbf{h}^{b}_{t}=\text{LSTM}(\mathbf{h}^{b}_{t+1},\mathbf{c}_{t}),$ (6)
where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward LSTM and backward LSTM respectively and $\mathbf{c}_{t}$ is the input at time step $t$. The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward LSTM and backward LSTM as $\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}]$. Please refer to the textbook [19] for a detailed description. Our BiLSTM module utilizes both past and future information and is capable of capturing long-range temporal relationships in both forward and backward directions. As empirically observed in Sec. 5.3, BiLSTM outperforms various alternatives, including average pooling, linear layer, traditional LSTM, and transformer [44]. Our final features output by BiLSTM not only are time-sensitive but also contain both local fine-grained and global semantic information, which are important for temporal reasoning in long videos. Unlike LongVLM [50], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and achieves better results as shown in Sec. 5.
The video tokens output by our BiLSTM module have dimensions of $(C*N^{c}_{f},~{}2*N_{V})$, while the LLM requires an input of dimensions $(N^{c}_{f},~{}N_{LLM})$. Here, $N_{V}$ and $N_{LLM}$ are the dimensions of the video tokens and the LLM latent space respectively, and $2$ is to account for the outputs of both forward and backward passes of BiLSTM. Thus, we pass the video tokens output by BiLSTM through projection layers to match the input dimensions required by the LLM.
3.3 Large Language Model
The LLM (i.e., yellow block in Fig. 2) takes as input the video tokens $\mathbf{X}^{v}$, query tokens $\mathbf{X}^{q}$, and generates responses $\mathbf{X}^{r}$ to users. Video LLMs typically employ a two-stage training scheme: i) leveraging large-scale image/video-text pairs for vision-language alignment to pre-train the model, and ii) utilizing instruction data for instruction following to fine-tune the pre-trained model. Following TimeChat [36], we use the checkpoint of the open-source LLaMA-2 7B model [43] and perform instruction tuning only. We optimize the below loss function:
$\mathcal{L}=-\log P_{\mathbf{\psi}}(\mathbf{X}^{r}|\mathbf{X}^{v},\mathbf{X}^{q})$ (7)
$=-\sum^{L}_{i=1}\log P_{\mathbf{\psi}}(x_{i}|\mathbf{X}^{v},\mathbf{X}^{q},\mathbf{X}^{r}_{<i}).$ (8)
Here, $\mathbf{\psi}$ is the learnable parameters of TemporalVLM, $L$ is the response length, $x_{i}$ is the current predicted token, and $\mathbf{X}^{r}_{<i}$ is the prior tokens appearing before $x_{i}$ in the response. Also, we employ LoRA [12] for efficient fine-tuning.
4 IndustryASM: A Large-Scale Long Video Dataset of Industry Assembly Processes
Figure 3: Example videos with different camera viewpoints, actors, backgrounds, and activities from our IndustryASM dataset.
We introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion studies.
4.1 Dataset Statistics
Our IndustryASM dataset includes 4803 videos with each video lasting 119 seconds on average, yielding 158 hours as the total dataset duration. In addition, it consists of 52 various industry assembly procedures with each assembly procedure having 12 steps on average. Example videos are shown in Fig. 3. Due to space limits, we provide additional details of IndustryASM in our supplementary material.
4.2 Annotation Process
Our industry assembly videos are manually annotated by industrial engineers into relevant framewise action segments. The action labels contain the same step naming conventions as used on factory floors during manufacturing processes. To ensure quality, each video is annotated by two labelers, i.e., one labeler obtains the labels for the video, while another labeler checks the labels. In cases of conflicts, both labelers discuss to address them. Overall, around 8% of the videos have conflicts and require fixing, yielding an agreement rate of roughly 92% between labelers. Given the framewise annotations we first convert them to timestamps by using the frame rate of the video. We manually write the instructions designed for the action segmentation task to ensure quality and generate the answers by using timestamps and action labels. Example instructions and responses are shown in Fig. S1 of our supplementary material.
5 Experiments
Tasks and Datasets. Our TemporalVLM model is evaluated on a few temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. For fair comparisons with TimeChat [36], we fine-tune our model on a subset [^1] We could not download the YT-Temporal [55] dataset on time due to its large size and the restricted number of downloads by YouTube. of the TimeIT [36] and Valley [28] datasets. The combined dataset contains a total of 142K long videos with diverse activities. We follow TimeChat [36] to use 6 instructions (generated by GPT-4 [1]) per task and convert the video annotations into chat samples. We evaluate on YouCook2 [59] for dense video captioning, Charades-STA [8] for temporal video grounding, QVHighlights [20] for video highlight detection, and our IndustryASM dataset for temporal action segmentation.
Implementation Details. We follow TimeChat [36] to use LLaMA-2 7B [43] as the LLM, and ViT-G/14 from EVA-CLIP [40] as the image encoder. We take InstructBLIP’s checkpoints to initialize the image Q-Former, and Video-LLaMA’s [57] checkpoints to initialize the video Q-Former. We employ pytorch’s [15] implementation of multiheaded attention in the fusion module of the clip encoder with 8 attention heads. The BiLSTM module includes a forward LSTM with one hidden layer and a backward LSTM with one hidden layer. The parameters of the BiLSTM module and projection layers are initialized randomly. The LLM is fine-tuned using LoRA [12] with rank 32. We use 6 clips and sample 96 frames from each clip. We fine-tune our model on a subset of TimeIT [36] and Valley [28] on a single 8-A100 (80 GB) machine. The parameters of the LLM and image encoder remain frozen throughout fine-tuning. Please see our supplementary material for more details.
Competing Methods. We compare our model against state-of-the-art video LLMs, including VideoChat-Embed [22], Video-LLaMA [57], Valley [28], TimeChat [36], and LongVLM [50]. Since LongVLM [50] is not trained on TimeIT [36] and Valley [28] and it is a multi-model pipeline (TimeChat [36] and TemporalVLM are end-to-end models), we replace our time-aware clip encoder and BiLSTM modules with their local and global feature extraction modules respectively while keeping the same image encoder and LLM, yielding our reimplemented LongVLM model which we use for evaluation. Nevertheless, we compare with the original LongVLM model [50] on general video understanding in our supplementary material.
Evaluation Metrics. We utilize different sets of metrics for evaluating temporal reasoning tasks:
  • Dense video captioning: The overall story-level evaluation is done through SODA_c [7]. The event localization performance is measured using F1 Score. CIDEr [45] is used to measure the quality of the generated captions.
  • Temporal video grounding: To gauge the effectiveness of temporal video grounding, we use R@1(IoU = $\alpha$), which is the percentage of predictions with an intersection over union greater than $\alpha$ compared to the ground truth.
  • Video highlight detection: We use mean average precision (mAP) and HIT@1 at IoU= 0.75 to evaluate the accuracy of the saliency scores for key moments in videos.
  • Temporal action segmentation: We report the framewise accuracy, F1 scores at overlapping thresholds $\{0.1,0.25,0.5\}$, and Edit distance. We use the intersection over union ratio to find the overlapping thresholds.
5.1 Zero-Shot Results
Model Dense Video Captioning Video Highlight Detection Temporal Video Grounding
YouCook2 [59] QVHighlights [20] Charades-STA [8]
SODA_c CIDEr F1 mAP HIT@1 R@1(IoU=0.5) R@1(IOU=0.7)
Valley [28] 0.1 0.0 1.5 10.9 15.2 4.7 1.6
Video-LLaMA [57] 0.0 0.0 0.1 11.3 15.6 2.7 1.2
VideoChat-Embed [22] 0.2 0.6 3.4 13.1 18.1 3.2 1.4
LongVLM [50] 0.8 2.5 12.3 11.1 15.0 13.9 6.1
TimeChat [36] 1.2 3.4 12.6 14.5 23.9 27.9 12.3
TemporalVLM (Ours) 1.2 3.7 13.1 16.4 31.3 30.1 13.2
Table 1: Zero-short results. We evaluate all methods on YouCook2 [59] for dense video captioning, QVHighlights [20] for video highlight detection, and Charades-STA [8] for temporal video grounding. Best results are in bold, while second best ones are underlined.
Dense Video Captioning Results. Tab. 1 shows that our TemporalVLM model achieves the best results across all three metrics, i.e., SODA_c, CIDEr, and F1 Score. For example, we outperform TimeChat [36] by +0.3 on CIDEr and +0.5 on F1 Score. This is likely because TimeChat [36] treats the entire video as a single clip and aggregates video tokens via query aggregation, overlooking fine-grained information. In contrast, our extracted features, which capture both local fine-grained cues from each clip and global temporal dependencies across multiple clips, are effective for caption generation and event localization. Moreover, our approach obtains superior results than LongVLM [50] across SODA_c (+0.4), CIDEr (+1.2), and F1 Score (+0.8), which confirms the importance of our time-aware clip encoder and BiLSTM for dense video captioning.
Video Highlight Detection Results. The video highlight detection results in Tab. 1 show that our TemporalVLM model obtains the best performance on both mAP and HIT@1. In particular, it surpasses TimeChat [36] by +1.9 in mAP and +7.4 in HIT@1, while outperforming LongVLM [50] by +5.3 at mAP and by +16.3 at HIT@1. These results suggest the effectiveness of our model for video highlight detection. By aggregating global features from time-aware local features, it identifies the most relevant timestamps and assigns them saliency scores based on their ranks of importance within the video against a given query. Though LongVLM [50] extracts multi-level features, the lack of a time-aware encoder hinders their results.
Temporal Video Grounding Results. It is evident from Tab. 1 that TemporalVLM achieves the best performance for temporal video grounding across R@1 (IoU=0.5) and R@1 (IoU=0.7). Specifically, it outperforms TimeChat [36] by +2.2 and +0.9 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively, and LongVLM [50] by +16.2 and +7.1 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively. The results validate the benefits of our time-aware and multi-level features for temporal video grounding. Our clip encoder extracts fine-grained cues for temporal reasoning within each clip via fusion, which enables our BiLSTM module to capture long-range temporal relationships between multiple clips to detect the given event effectively. The lack of a time-aware encoder affects LongVLM [50], while TimeChat [36] suffers from a lack of fine-grained information.
5.2 Supervised Results
Dense Video Captioning, Video Highlight Detection, and Temporal Video Grounding Results. We present the results of all methods in the supervised setting in Tab. 2. Our TemporalVLM model achieves the best results across all tasks and metrics. As compared to the zero-shot results in Tab. 1, we see a further increase in performance, e.g., +24.3 on R@1 (IoU=0.5) in the temporal video grounding task and +9.5 on CIDEr in the dense video captioning task. However, we acknowledge that task-specific models such as MMN [49] and QD-DETR [30] may outperform generalist models such as TimeChat [36] and our TemporalVLM model due to their task-specific design paradigms, e.g., using a mutually matching network to directly model the relationship between the query text and video content or specialized tokens and loss functions during fine-tuning.
Model Dense Video Captioning Video Highlight Detection Temporal Video Grounding
YouCook2 [59] QVHighlights [20] Charades-STA [8]
SODA_c CIDEr F1 mAP HIT@1 R@1 (IoU=0.5) R@1 (IoU=0.7)
LongVLM [50] 2.3 8.1 16.9 16.0 22.5 27.2 11.9
TimeChat [36] 3.1 10.3 19.5 21.7 37.9 46.7 23.7
TemporalVLM (Ours) 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 2: Supervised results. We first fine-tune all models on TimeIT [36] and Valley [28], and then perform task-specific fine-tuning and evaluation on YouCook2 [59] for dense video captioning, QVHighlights [20] for video highlight detection, and Charades-STA [8] for temporal video grounding. Best results are in bold, while second best ones are underlined.
Temporal Action Segmentation Results. We further test all methods on our IndustryASM dataset for temporal action segmentation. Since the activities in IndustryASM significantly differ from those in TimeIT [36] and Valley [28], we fine-tune all models on IndustryASM before evaluation. Tab. 3 presents the results. TemporalVLM performs the best across all metrics. For example, it outperforms TimeChat [36] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [50] by +8.9 and +5.3 on Edit and Acc respectively. The results confirm the benefits of our fused time-aware and multi-level features.
Model F1@{10,25,50} Edit Acc
TimeChat [36] {11.1, 8.5, 4.4} 32.7 46.0
LongVLM [50] {17.1, 13.3, 7.3} 47.3 47.6
TemporalVLM (Ours) {22.3, 18.3, 11.1} 56.2 52.9
Table 3: Temporal action segmentation results in the supervised setting on IndustryASM. All models are fine-tuned on IndustryASM before evaluation. Best results are in bold, while second best ones are underlined.
Multi-Clip Encoder vs. Time-Aware Encoder. Multi-clip encoders in LongVLM [50] and TemporalVLM are crucial to capturing fine-grained cues, leading to superior results over TimeChat [36] on temporal action segmentation in Tab. 3. In contrast, time-aware encoders in TimeChat [36] and TemporalVLM are important to event localization and temporal modeling, resulting in better performance than LongVLM [50] on dense video captioning, video highlight detection, and temporal video grounding in Tabs. 1 and 2. Our TemporalVLM encoder is both multi-clip and time-aware, yielding the best performance across all tasks.
5.3 Ablation Results
Below we study the effects of our overlapping sliding video Q-Former and BiLSTM. For this purpose, we fine-tune multiple configurations of TemporalVLM on YouCook2 [59].
Impacts of Overlapping Sliding Video Q-Former. We conduct experiments in Tab. 4 to demonstrate the importance of fusing encoded frames across overlapping temporal windows in our clip encoder. Our model with fusion across overlapping windows achieves the best results, whereas using non-overlapping windows or removing the fusion module leads to performance drops. The results validate the design of our overlapping sliding video Q-Former module.
Model SODA CIDEr F1
No Overlap + No Fusion 2.3 11.0 17.3
No Overlap + Fusion 2.0 9.5 16.2
Overlap + No Fusion 2.8 11.7 18.9
Overlap + Fusion 3.4 13.2 20.0
Table 4: Effects of overlapping sliding video Q-Former. Best results are in bold, second best are underlined.
Impacts of BiLSTM. To study the effects of BiLSTM, we replace it with various alternatives and report the results in Tab. 5. From Tab. 5, BiLSTM outperforms average pooling, which is not trained and employed in prior works [29, 50], and traditional LSTM, which only performs the forward pass and hence relies only on past cues. In contrast, BiLSTM is learnable and conducts both forward and backward passes to utilize both past and future cues. Moreover, we replace BiLSTM with linear layer to examine the impacts of video Q-Formers [57, 36] alone, leading to worse results. This highlights the importance of BiLSTM even when video Q-Formers [57, 36] are being used. Finally, BiLSTM outperforms transformer [44], which we will discuss below.
Model # Layers SODA_c CIDEr F1
Average Pooling 1 1.7 6.0 16.1
Linear Layer 1 2.4 9.5 17.8
Transformer 2 2.3 6.8 15.6
LSTM 1 2.8 9.4 18.2
BiLSTM 2 3.4 13.2 20.0
Table 5: Effects of the BiLSTM module. Best results are in bold, while second best ones are underlined.
(Bi)LSTM vs. Transformer. The permutation-invariant nature of the self-attention mechanism in transformer [44] leads to loss in temporal information of the input sequence. Though using positional encoding alleviates this issue, it is inevitable that some temporal information loss occurs [56]. In contrast, (Bi)LSTM processes data one by one and preserves the order of the sequence, making them better suited for temporal modeling. Similar to [56], we also observe in Tab. 5 that linear layer outperforms transformer [44].
5.4 Qualitative Results
Figure 4: Example dense video captioning results in the zero-shot setting on a YouCook2 [59] video. Inaccuracies are shown in red.
Figure 5: Example temporal action segmentation results in the supervised setting on an IndustryASM video. Colored segments represent predicted actions with a particular color denoting the same action across all models. Black represents background frames.
Dense Video Captioning Results. Fig. 4 shows example results of dense video captioning in the zero-shot setting on a YouCook2 [59] video. TimeChat [36] fails to predict any correct timestamps and tends to hallucinate. LongVLM [50] predicts some timestamps correctly and generates more accurate captions, although it still hallucinates objects not present in the video and produces inaccurate/repeated captions. In contrast, TemporalVLM produces much more accurate timestamps and captions. Also, it is the only model that provides captions till the end of the video, showing the effectiveness of the model for long videos.
Temporal Action Segmentation Results. Fig. 5 illustrates example results of temporal action segmentation in the supervised setting on an IndustryASM video. TimeChat [36] falls short in segmenting the video correctly and hallucinates actions not present in the video. LongVLM [50] fares better in terms of the actions detected in the video but struggles with predicting the action segments and hallucinations further into the video. In contrast, TemporalVLM predicts action segments significantly closer to the ground truth.
6 Conclusion
We propose TemporalVLM, a video LLM for effective temporal reasoning and fine-grained understanding in long videos. Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts fused time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features. The extracted features not only are time-sensitive but also contain both local and global cues. In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for time and motion studies and temporal action segmentation evaluation. Lastly, we conduct extensive experiments on TimeIT [36] and IndustryASM to demonstrate our superior results over previous methods across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to incorporate LSTMs into video LLMs.
Appendix A Supplementary Material
In this supplementary material, we first provide additional details of our IndustryASM dataset and our TemporalVLM implementation in Secs. A.1 and A.4 respectively. Next, we provide a comparison with specialized models such as Vid2Seq [53] and QD-DETR [30] in Sec. A.5, followed by demonstrating the effects of number of training epochs in Sec. A.6. Sec. A.7 compares the sizes of our model, TimeChat [36], and LongVLM [50], while Sec. A.8 assesses the impact of number of clips. Sec. A.9 evaluates our model for general video understanding. We then present several qualitative results in Sec. A.10, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results. Finally, we discuss the limitations and societal impacts of our work in Sec. A.11.
A.1 IndustryASM Details
A.2 Dataset Statistics
Our IndustryASM dataset comprises of 4803 videos in total and the average video duration is 119 seconds. Therefore, the total dataset duration is 158 hours. These videos are distributed among 52 industry assembly processes or datasets, ranging from automotive manufacturing, electronic device manufacturing, medical device manufacturing to heating, ventilation, and air conditioning (HVAC) manufacturing. Each assembly process or dataset involves 12 steps or actions on average. We categorize each step into one of the following classes:
  • Moving: In this category, the worker(s) transfer parts or subassemblies within or between workstations, relocating items from one position to another within the workspace.
  • Assembling: Here, the worker(s) actively build or complete subassemblies by adding components or combining parts to form a complete unit.
  • Positioning: This involves the worker(s) placing parts onto a subassembly or mounting a subassembly in a designated position within the workstation.
  • Packaging: In this category, the worker(s) place subassemblies or components into boxes or other forms of packaging materials.
Examples of these 4 categories are presented in Fig. 3 of the main paper. We further classify each assembly process or dataset into one category which appears the most among all of its steps. For example, the Air_Cleaner dataset includes 10 steps (i.e., 4 moving steps, 3 assembling steps, and 3 positioning steps), and hence it is classified as “moving” which appears the most among its 10 steps. We provide the statistics for each category as well as the entire IndustryASM dataset in Tab. 6.
(a) Complete Dataset
Number of videos 4803
Average video duration 119 seconds
Total dataset duration 158 hours
Average number of steps 12
Number of datasets 52
(b) Moving
Number of videos 383
Average video duration 176 seconds
Total dataset duration 19 hours
Average number of steps 11
Number of datasets 7
(c) Assembling
Number of videos 1671
Average video duration 144 seconds
Total dataset duration 67 hours
Average number of steps 14
Number of datasets 20
(d) Positioning
Number of videos 1696
Average video duration 109 seconds
Total dataset duration 51 hours
Average number of steps 13
Number of datasets 18
(e) Packaging
Number of videos 1053
Average video duration 74 seconds
Total dataset duration 22 hours
Average number of steps 10
Number of datasets 7
Table 6: Dataset statistics for IndustryASM.
A.3 Annotation Process
Our IndustryASM videos are annotated with action names and action timestamps by industrial engineers. To ensure quality, two labelers are assigned for each video, i.e., one labeler provides the labels for the video, while another labeler checks the labels. If there are conflicts, both labelers discuss to fix them. In general, roughly 8% of the videos have conflicts and need fixing, resulting in an agreement rate of around 92% between labelers. We manually write the user instructions for the temporal action segmentation task and generate the ground truth responses by using the action names and action timestamps. Fig. 6 illustrates examples of our instruction and ground truth response for the temporal action segmentation on an IndustryASM video.
Figure 6: Examples of our instruction and target response for the temporal action segmentation task on an IndustryASM video.
A.4 Implementation Details
We present a detailed list of hyperparameter values used in tuning our TemporalVLM model in Tab. 7. We sample $C=6$ clips from the video, with each consisting of $N^{c}_{f}=96$ frames. The frame encoder processes frames of sizes $224\times 224$ across $14\times 14$ patches. Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively. Inside the clip encoder, we divide the timestamp encoded input to the video Q-former into windows of size $q$ and overlap $o$. We then fuse the concatenated window-wise outputs of the Q-former via multiheaded self attention with 8 attention heads that have the same hidden size $N_{v}$ as the Q-Former. The output of the clip encoder contains a fused time-aware representation of each clip. We then concatenate the time-aware clip features according to their temporal order of appearance in the video. This sequence is of dimensions $(C*N^{c}_{f},~{}N_{V})$, with $N_{V}$ denoting the dimension of the video tokens. This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C*N^{c}_{f},~{}2*N_{V})$. The BiLSTM module comprises of two hidden layers, with one hidden layer for the forward LSTM and another hidden layer for the backward LSTM. Projection layers are used to project the BiLSTM output into the LLM input space. The first layer projects the BiLSTM output to the dimensions of $(C*N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM.
Hyperparameter Value
Number of clips $C$ 6
Number of frames sampled per clip $N^{c}_{f}$ 96
Frame encoder patch size 14$\times$14
Frame resolution 224$\times$224
Frame sampling type uniform
Number of epochs 7
Batch size 32
Learning rate 5e-5
Warmup learning rate 5e-6
Weight decay 0.01
Optimizer Adam
AdamW $\beta$ (0.9,0.999)
Scheduler LinearWarmupCosineLR
Q-Former window size $q$ 32
Overlap $o$ 16
Number of visual tokens per window 32
Number of layers in clip encoder 2
Number of layers in image encoder 12
No of attention heads in fusion module 8
Q-Former hidden size $N_{V}$ 768
Fusion module hidden size $D$ 768
BiLSTM input size 768
BiLSTM hidden size 768
LLM hidden size $N_{LLM}$ 4096
Table 7: Hyperparameter settings.
A.5 Comparison With Specialized Models
Tab. 8 shows generalist models underperform specialized models such as Vid2Seq  [53] and QD-DETr [30] whose advantages come from task-specific training data and model designs. In particular, the use of 4 loss functions for 200 epochs during training and the utilization of saliency tokens for saliency prediction in QD-DETR  [30] leads to superior performance in highlight detection tasks. In contrast our TemporalVLM is trained on a simpler language modeling loss on lesser epochs. Generalist models such as ours exhibit better generalization across zero-shot, multi-task, and multi-domain settings.
Model DVC VHD TVG
SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
Vid2Seq 7.9 47.1 27.3
QD-DETR 38.9 62.4
QD-DETR w/ Audio 39.0 62.9
MMN 50.5 29.7
VDI 52.3 31.4
TemporalVLM 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 8: Comparisons with specialized models.
A.6 Effects of Number of Training Epochs
Tab. 9 shows the performance of our models across different training epochs.Since our model uses a learnable fusion layer in the clip encoder and a BiLSTM module to aggregate clip level features into a global representation, we observe the best results by training for 7 epochs. However, Tab. 9 reports our results with 3 epochs, which underperform those with 7 epochs but still outperform those of TimeChat and LongVLM.
Model DVC VHD TVG
SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
LongVLM (3 epochs) 2.3 8.1 16.9 16.0 22.5 27.2 11.9
TimeChat (3 epochs) 3.1 10.3 19.5 21.7 37.9 46.7 23.7
TemporalVLM (3 epochs) 3.2 12.9 19.7 23.9 42.3 50.0 25.9
TemporalVLM (7 epochs) 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 9: Effects of number of training epochs.
A.7 Comparison on Model Sizes
Tab. 10 compares the sizes of our TemporalVLM model, TimeChat [36], and LongVLM [50] in terms of the number of learnable parameters. Our TemporalVLM model includes a trainable BiLSTM module and projection layers, leading to a 5% and 4% increase in the number of trainable parameters over TimeChat [36] and LongVLM [50] respectively. Nevertheless, our TemporalVLM model achieves the best performance across various temporal reasoning and fine-grained understanding tasks, despite using less training data than TimeChat [36]. In particular, TimeChat [36] is trained on the complete TimeIT [36] and Valley [28] datasets, whereas our TemporalVLM model is trained on a subset [^2] We could not download the YT-Temporal [55] dataset on time due to its large size and the restricted number of downloads by YouTube. of the TimeIT [36] and Valley [28] datasets.
A.8 Impacts of Number of Short-Term Clips
Tab. 11 presents the effects of the number of short-term clips on our TemporalVLM model. Due to memory limitations, we have experimented with three values for the number of short-term clips, namely 2, 4, and 6. It is evident from Tab. 11 that the performance is improved with increasing the number of short-term clips, since our model is able to access more data from the input video.
Model Number of Trainable Parameters
TimeChat [36] 241 millions
LongVLM [50] 244 millions
TemporalVLM (Ours) 255 millions
Table 10: Comparison on model sizes.
# Clips SODA_c CIDEr F1
2 2.6 9.2 18.5
4 3.1 11.3 18.9
6 3.4 13.2 20.0
Table 11: Effects of the number of short-term clips. Best results are in bold, while second best ones are underlined.
A.9 General Video Understanding Results
Method Data CI DO CU TU C
Video-LLaMA [57] 10M 1.96 2.18 2.16 1.82 1.79
Video-ChatGPT [29] 100k 2.50 2.57 2.69 2.16 2.20
Valley [28] 234k 2.43 2.13 2.86 2.04 2.45
BT-Adapter [25] 10M 2.68 2.69 3.27 2.34 2.46
LongVLM [50] 100k 2.76 2.86 3.34 2.39 3.11
TemporalVLM 100k 2.88 2.91 3.45 2.50 3.16
Table 12: General video understanding results. Best results are in bold, while second best ones are underlined.
Following LongVLM [50], we evaluate the performance of our TemporalVLM model on the general video understanding benchmark provided by Video-ChatGPT [29] in Tab. 12. The evaluation metrics include Correctness Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (C). It is evident from Tab. 12 that our TemporalVLM model achieves the best performance across all metrics, outperforming all competing methods, including the original LongVLM model [50].
A.10 Qualitative Results
In addition to Figs. 4 and 5 in the main paper, we provide additional qualitative results in this section. In particular, Figs. 7, 8, and 9 present qualitative results by TemporalVLM in the supervised setting for dense video captioning, video highlight detection, and temporal video grounding respectively. More importantly, qualitative results demonstrating the generalization abilities of TemporalVLM in the zero-shot setting are shown in Fig. 10. Overall, our TemporalVLM model demonstrates promising performance in a variety of temporal reasoning and fine-grained understaning tasks.
Figure 7: Qualitative examples highlighting the dense video captioning capabilities of TemporalVLM in the supervised setting. The model is asked to provide the timestamps of actions that occur in the videos along with brief descriptions of the actions.
Figure 8: Qualitative examples demonstrating the video highlight detection capabilities of TemporalVLM in the supervised setting. The model is given a video along with an action description. It is prompted to provide the frames that match the action description and their saliency scores.
Figure 9: Qualitative examples showing the temporal video grounding capabilities of TemporalVLM in the supervised setting. A video and a query is given to the model. It is prompted to provide the timestamps at which the query occurs. Magenta represents the ground truth and predicted timestamps for the first query, while Orange indicates those for the second query.
Figure 10: Examples of the generalization capabilities of TemporalVLM in the zero-shot setting. (Top) The model is prompted to provide the timestamps of actions and brief descriptions of actions for an egocentric video of furniture assembly. (Bottom) The model is provided with a documentary film. It is asked to predict the timestamps when the query happens. Blue denotes the ground truth and predicted timestamps.
A.11 Discussions
A.12 Limitations
Despite promising performance in several temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation, our time-aware video LLM may struggle with complex/dense temporal reasoning tasks and videos with extreme durations. Our future works will enhance TemporalVLM for tackling complex/dense temporal reasoning tasks and videos with extreme durations.
A.13 Societal Impacts
Our time-sensitive video LLM could enable a variety of applications including optimizing and tracking industry assembly processes. More specifically, for assembly process optimization, industrial engineers could utilize TemporalVLM to automatically decompose a video recording of an assembly process into segments as well as generate a brief description for each segment. In addition, TemporalVLM could improve the performance of frontline workers by tracking their assembly process and notify them as soon as they miss a step. Nevertheless, we acknowledge that our time-aware video LLM could be misused for surveillance and monitoring of individuals, which emphasizes the importance of responsible AI principles to guide the use of this technology.